In the right triangle, the ratio of the measures of the two acute angles is 4:1. What is the measure in degrees, of the larger acute angle?
Answer by stanbon(75887)

In the right triangle, the ratio of the measures of the two acute angles is 4:1. What is the measure in degrees, of the larger acute angle?
------------------------------------
4:1 is the same as 4x:x
-----------------
EQUATION:
4x + x = 90
5x = 90
x = 18
-------------------------
The larger angle is 4x = 4*18 = 72 degrees
===========================
Cheers,
Stan H.